OpenStudy (ledah):

Simplify the following: 5 x^3 y +x y^4+ 2xy divided by xy

OpenStudy (ledah):

oops that should be changed quite a bit.. 5x^3 y +xy^4+2xy

OpenStudy (ledah):

the spaces were too far apart..

OpenStudy (unklerhaukus):

\[\frac{5 x^3 y +x y^4+ 2xy }{xy}=\frac{xy(5 x^2 + y^3+ 2) }{xy}\]

OpenStudy (apoorvk):

take 'xy' out from each term, and then take a common. The xy in the numerator, would get cancelled by the 'xy' by which you need to divide the expression.

OpenStudy (ledah):

cancel out the "xy"? with each match to it?

OpenStudy (apoorvk):

yeah, if 'match' means this: \[\cancel{xy}\]

OpenStudy (ledah):

thats what I mean o_o

OpenStudy (unklerhaukus):

\[\frac{5 x^3 y +x y^4+ 2xy }{xy}=\frac{\cancel{xy}(5 x^2 + y^3+ 2) }{\cancel {xy}}=5 x^2 + y^3+ 2\]

OpenStudy (ledah):

yeah I got that. thanks you guys